# Stochastic Optimal Control of an Industrial Power-to-Heat System with High-Temperature Heat Pump and Thermal Energy Storage

### v2

@@ Appendix E Proofs
As for the proof above we use the definitions of $I_{W}(t),I_{S,1}(t)$ and $I_{S,2}(t)$. Notice that all of these expressions are martingales with mean zero. Also note that $B^{W}$ and $B^{S}$ are independent Brownian motions and therefore $I_{W}(t)$ is independent of $I_{S,1}(t)$ and $I_{S,2}(t)$. Using the solutions for $Y^{W}(t)$ and $Y^{S}(t)$ in (91) and (103), respectively, leads to the following calculation of $\mathbb{E}[Y^{W}(t)Y^{S}(t)]$
$\mathbb{E}[Y^{W}(t)Y^{S}(t)]$ $=\mathbb{E}[(m_{Y^{W}}(\tau,y_{W})+I_{W}(t))(m_{Y^{S}}(\tau,y_{W},y_{S})+I_{S,1}(t)+I_{S,2}(t))]$ (122)
$=\mathbb{E}[m_{Y^{W}}(\tau,y_{W})m_{Y^{S}}(\tau,y_{W},y_{S})]+\mathbb{E}[m_{Y^{W}}(\tau,y_{W})I_{S,1}(t)]+\mathbb{E}[m_{Y^{W}}(\tau,y_{W})I_{S,2}(t)]$ (123)
$~{}~{}~{}+\mathbb{E}[I_{W}(t)m_{Y^{S}}(\tau,y_{W},y_{S})]+\mathbb{E}[I_{W}(t)I_{S,1}(t)]+\mathbb{E}[I_{W}(t)I_{S,2}(t)]$ (124)
$=m_{Y^{W}}(\tau,y_{W})m_{Y^{S}}(\tau,y_{W},y_{S})+\mathbb{E}[I_{W}(t)]\mathbb{E}[I_{S,1}(t)]+\mathbb{E}[I_{W}(t)I_{S,2}(t)]$ (125)
@@
Substitution into (121) and applying Itô isometry yields for the covariance $\Sigma_{WS}(\tau)$ the following result
$\Sigma_{WS}(t)$ $=\mathbb{E}[I_{W}(t)I_{S,2}(t)]$ (127)
$=\mathbb{E}\bigg{[}\bigg{(}\int_{t_{a}}^{t}\sigma_{W}e^{-\lambda_{W}(t-r)}\mathrm{d}B^{W}(r)\bigg{)}\bigg{(}-\tfrac{\lambda_{S}c_{W}}{\lambda_{S}-\lambda_{W}}\sigma_{W}\int_{t_{a}}^{t}(e^{-\lambda_{W}(t-r)}-e^{-\lambda_{S}(t-r)})\mathrm{d}B^{W}(r)\bigg{)}\bigg{]}$ (128)
$=\int_{t_{a}}^{t}-\tfrac{\lambda_{S}c_{W}}{\lambda_{S}-\lambda_{W}}\sigma_{W}^{2}e^{-\lambda_{W}(t-r)}(e^{-\lambda_{W}(t-r)}-e^{-\lambda_{S}(t-r)})\mathrm{d}r$ (129)
$=-\tfrac{\lambda_{S}c_{W}}{\lambda_{S}-\lambda_{W}}\sigma_{W}^{2}\int_{t_{a}}^{t}(e^{-2\lambda_{W}(t-r)}-e^{-(\lambda_{S}+\lambda_{W})(t-r)})\mathrm{d}r$ (130)
$=-\tfrac{\lambda_{S}c_{W}}{\lambda_{S}-\lambda_{W}}\sigma_{W}^{2}\bigg{(}\int_{t_{a}}^{t}e^{-2\lambda_{W}(t-r)}\mathrm{d}r-\int_{t_{a}}^{t}e^{-(\lambda_{S}+\lambda_{W})(t-r)}\mathrm{d}r\bigg{)}$ (131)
$=-\tfrac{\lambda_{S}c_{W}}{\lambda_{S}-\lambda_{W}}\sigma_{W}^{2}\bigg{(}\tfrac{\Sigma_{Y^{W}}^{2}(\tau)}{\sigma_{W}^{2}}-\tfrac{1}{(\lambda_{S}+\lambda_{W})}\bigg{(}1-e^{-(\lambda_{S}+\lambda_{W})\tau}\bigg{)}\bigg{)}$ (132)
@@
where $\int_{t_{a}}^{t}e^{-2\lambda_{W}(t-r)}\mathrm{d}r=\Sigma_{Y^{W}}^{2}(\tau)/\sigma_{W}^{2}$ in the second to last equality.
In order to show that $\Sigma_{WS}(\tau)\leq 0$ one has to consider two different cases.
@@
In this case it holds that $\lambda_{S}-\lambda_{W}>0$ and therefore $-\tfrac{\lambda_{S}c_{W}}{\lambda_{S}-\lambda_{W}}\leq 0.$ Also note that $\lambda_{S}+\lambda_{W}>2\lambda_{W}$ and using (93) is leading to
@@
Now it holds that $\lambda_{S}-\lambda_{W}<0$ giving $-\tfrac{\lambda_{S}c_{W}}{\lambda_{S}-\lambda_{W}}>0.$ Conversely to case 1 we get that $\lambda_{S}+\lambda_{W}<2\lambda_{W}$ leading to
@@
being zero if and only if $c_{W}=0$ which completes the proof.
@@
The matrix $\Sigma(\tau)$ is positive definite if $\operatorname{\mathrm{tr}\,}\Sigma(\tau)>0$ and $\det\Sigma(\tau)>0$. Since $\Sigma_{W}^{2}(\tau),\Sigma_{S}^{2}(\tau)>0$ we get that $\operatorname{\mathrm{tr}\,}\Sigma(\tau)=\Sigma_{W}^{2}(\tau)+\Sigma_{S}^{2}(\tau)>0$.
On the other hand, $\det\Sigma(\tau)=\Sigma_{W}^{2}(\tau)\Sigma_{S}^{2}(\tau)-\Sigma_{WS}^{2}>0$ if $\Sigma_{W}^{2}(\tau)\Sigma_{S}^{2}(\tau)>\Sigma_{WS}^{2}(\tau)$. Now let $\gamma=\bigg{(}\tfrac{\lambda_{S}c_{W}\sigma_{W}}{\lambda_{S}-\lambda_{W}}\bigg{)}^{2}$, then calculations yield
$\Sigma_{W}^{2}(\tau)\Sigma_{S}^{2}(\tau)$ $=\Sigma_{W}^{2}(\tau)\bigg{[}\Sigma_{Y^{S}}^{2}(\tau)+\gamma\big{(}\Sigma_{Y^{W}}^{2}(\tau)+\frac{\sigma_{W}^{2}}{\sigma_{S}^{2}}\Sigma_{Y^{S}}^{2}(\tau)-\frac{2\sigma_{W}^{2}}{\lambda_{S}+\lambda_{W}}\big{(}1-e^{-(\lambda_{S}+\lambda_{W})\tau}\big{)}\big{)}\bigg{]}$ (138)
$=\Sigma_{W}^{2}(\tau)\Sigma_{Y^{S}}^{2}(\tau)+\gamma\bigg{(}\Sigma_{Y^{W}}^{4}(\tau)+\frac{\sigma_{W}^{2}}{\sigma_{S}^{2}}\Sigma_{W}^{2}(\tau)\Sigma_{Y^{S}}^{2}(\tau)-\frac{2\Sigma_{W}^{2}(\tau)\sigma_{W}^{2}}{\lambda_{S}+\lambda_{W}}\big{(}1-e^{-(\lambda_{S}+\lambda_{W})\tau}\big{)}\bigg{)}$ (139)
$=\Sigma_{W}^{2}(\tau)\Sigma_{Y^{S}}^{2}(\tau)+\gamma\bigg{(}\bigg{[}\Sigma_{W}^{2}(\tau)-\frac{\sigma_{W}^{2}}{\lambda_{S}+\lambda_{W}}(1-e^{-(\lambda_{S}+\lambda_{W})\tau})\bigg{]}^{2}$ (140)
$\hskip 142.26378pt+\frac{\sigma_{W}^{2}}{\sigma_{S}^{2}}\Sigma_{W}^{2}(\tau)\Sigma_{Y^{S}}^{2}(\tau)-\frac{\sigma_{W}^{4}}{(\lambda_{S}+\lambda_{W})^{2}}(1-e^{-(\lambda_{S}+\lambda_{W})\tau})^{2}\bigg{)}$ (141)
@@
where $A=\frac{\sigma_{W}^{2}}{\sigma_{S}^{2}}\Sigma_{W}^{2}(\tau)\Sigma_{Y^{S}}^{2}(\tau)-\frac{\sigma_{W}^{4}}{(\lambda_{S}+\lambda_{W})^{2}}(1-e^{-(\lambda_{S}+\lambda_{W})\tau})^{2}$. Note that $\Sigma_{W}^{2}(\tau)\Sigma_{Y^{S}}^{2}(\tau)>0$ and $\gamma>0$, thus we need to verify that $A\geq 0$ in order to show that $\Sigma$ is positive definite. Using (93) and (93) we obtain
@@
and that $\int_{0}^{1}e^{-az}\,\mathrm{d}z=\frac{1-e^{-a}}{a}$ leading to
@@
Applying the Cauchy-Schwartz inequality yields
@@
E.2 Proof of Proposition 3.2
Let $0\leq t_{a}<t_{b}\leq t_{\text{E}}$ with $t_{a}=t_{n}$ and $t_{b}=t_{n+1}$ and $(Y^{W}_{n},Y^{S}_{n})=(y_{W},y_{S})$. The pair $(\log W_{n+1},S_{n+1})$ is bivariate Gaussian since it results from a linear transformation of the bivariate Gaussian $(Y^{W}_{n+1},Y^{S}_{n+1})$. Moreover, we have that $\Delta t=\tau=t_{n+1}-t_{n}=const$ and therefore the covariance matrix is given by $\Sigma=\Sigma_{Y}(\Delta t)=\Sigma_{Y}(\tau)$. The linear transformation also results in the following mean
@@
$m^{W}_{n+1}(w)$ $=\mu_{W}(t_{n+1})+m_{Y^{W}}(\Delta t,y_{W}),$ (147)
$m^{S}_{n+1}(w,s)$ $=\;\mu_{S}(t_{n+1})\,+m_{Y^{S}}(\Delta t,y_{W},y_{S}),$ (148)
@@
E.3 Proof of Corollary 3.3
The Cholesky decomposition for $\Sigma$ can be verified by some simple calculations. Furthermore, recalling Proposition 3.2, the pair $(\log W,S)$ is bivariate Gaussian with mean $m_{WS}$ and covariance matrix $\Sigma$. Therefore, given a standard normally distributed random vector $Z_{n+1}=(Z^{W}_{n+1},Z)^{S}_{n+1})^{\top}\sim\pazocal{N}(0_{2},I_{2})$ we can write $(\log W,S)$ as a linear transformation of the form
$(\log W_{n+1}S_{n+1})=m^{WS}_{n+1}(W_{n},S_{n})+AZ_{n+1}.$ (149)
@@
Appendix F Calibration of Wind Speed and Energy Price Model
@@
The wind speed and the energy price are modeled by continuous-time Ornstein-Uhlenbeck processes on $[0,t_{\text{E}}]$ see (16). Choosing appropriate parameters is important in order to obtain a realistic dynamical system. Therefore, real-world data are needed to calibrate the model parameters. Data for calibration can be fetched from the website of Deutscher Wetterdienst  [^1] Deutscher Wetterdienst, Climate data center, https://opendata.dwd.de/climate_environment and ENTSO-E  [^2] ENTSO-E, Transparency platform, https://transparency.entsoe.eu/dashboard/show. Note that the electricity price data is provided in a resolution of every 15 minutes or every hour. The latter results by averaging the 15 minute values. In our experiments wind speed and electricity price are provided in an hourly resolution for the entire year 2020, i.e. 8760 data points each. Note that the model assumes that data for wind speed is given by realizations of a random variable $e^{Z(t)}$ for a Gaussian process $Z$. Therefore, the fetched data has to be transformed by applying the logarithm. Denote the individual datasets by $z^{\dagger}_{1},\ldots,z^{\dagger}_{8760}$ with $\dagger\in\{W,S\}$, where $z_{j}^{\dagger}$ is the data for the $j$-th hour for wind and price respectively. Before doing estimations, outliers need to be removed from the dataset. This is realized by applying the $3$-$\sigma$ rule and remove all trajectories with values outside of the $99.7$ confidence interval.
F.1 Estimation of the Seasonality Function
@@
For an approximation of the seasonality functions $\mu_{W}$ and $\mu_{S}$ an appropriate modeling function is fitted to this data by least-square regression. The following models where chosen in order capture the daily, half-daily and yearly trends in the underlying data. For wind speed
$\mu_{W}(t)=k^{W}_{0}+k^{W}_{1}\cos\bigg{(}\frac{2\pi(t-t^{W}_{1})}{8760}\bigg{)}+k^{W}_{2}\cos\bigg{(}\frac{2\pi(t-t^{W}_{2})}{24}\bigg{)}$ (150)
and electricity price respectively
$\mu_{S}(t)=k^{S}_{0}+k^{S}_{1}\cos\bigg{(}\frac{2\pi(t-t^{S}_{1})}{8760}\bigg{)}+k^{S}_{2}\cos\bigg{(}\frac{2\pi(t-t^{S}_{2})}{24}\bigg{)}+k^{S}_{3}\cos\bigg{(}\frac{2\pi(t-t^{S}_{3})}{12}\bigg{)}.$ (151)
@@
The denominator in the cosine functions reflects the daily, half-daily and yearly components of the data, respectively. The choice of these functions is motivated by the patterns in the data. Wind speeds are commonly higher in winter compared to summer month. This fact is also reflected in the daily variations, where wind is slower at night and faster during the day. Electricity prices need to adjust to the changing weather conditions, especially with regard to renewable energy resulting in a yearly seasonality. This adjustments are also needed on a daily scale arising due to day-to-night cycle of renewable energy sources. The price data reveals that there is a need to include a half-daily trend, due to the fact that most electricity is used during two periods in a day. Firstly, in the morning when people wake up and in the afternoon when they return from work. The fitted parameters for the seasonalites can be found in Table 2.
@@
F.2 Estimation of Ornstein-Uhlenbeck Parameters
Recall that $\log W(t)=\mu_{W}(t)+Y^{W}(t)$ and $S(t)=\mu(t)+Y^{S}(t)$, where $Y^{W}$ and $Y^{S}$ are Ornstein-Uhlenbeck processes described in Lemma 3.1 in equation (17). The discrete-time analogue of the closed-form solution (18) is a autoregressive process of the form
$Y^{W}_{n+1}$ $=p_{W}Y_{n}^{W}+\Sigma_{Y^{W}}Z^{W}_{n+1}$ (152)
$Y^{S}_{n+1}$ $=p_{S}Y_{n}^{S}+q_{S}Y_{n}^{W}+\Sigma_{S}Z^{S}_{n+1}$ (153)
with a sequence $(Z_{n})_{n=1,\ldots,N}$ of independent standard normally distributed random vectors $Z_{n}=(Z_{n}^{W},Z_{n}^{S})^{\top}\sim\pazocal{N}(0_{2},I_{2})$. The parameters $p_{W},p_{S},q_{S},\Sigma_{Y^{W}}^{2}$ and $\Sigma_{S}^{2}$ are identified with their associated quantities given in equation (20), that is
$\begin{split}&p_{W}=e^{-\lambda_{W}\Delta t},\qquad\Sigma_{Y^{W}}^{2}=\tfrac{\sigma_{W}^{2}}{2\lambda}(1-e^{-2\lambda_{W}\Delta t})\\ &p_{S}=e^{-\lambda_{S}\Delta t},\qquad q_{S}=-\frac{\lambda_{S}c_{W}}{\lambda_{S}-\lambda_{W}}(e^{-\lambda_{W}\Delta t}-e^{-\lambda_{S}\Delta t}),\\ &\Sigma_{S}^{2}=\Sigma_{Y^{S}}^{2}+\tfrac{(\lambda_{S}c_{W})^{2}}{(\lambda_{S}-\lambda_{W})^{2}}\bigg{[}\Sigma_{Y^{W}}^{2}+\tfrac{\sigma_{W}^{2}}{\sigma_{S}^{2}}\Sigma_{Y^{S}}^{2}-\tfrac{2\sigma_{W}^{2}}{\lambda_{S}+\lambda_{W}}(1-e^{-(\lambda_{S}+\lambda_{W})\Delta t})\bigg{]}.\end{split}$ (154)
Let $z^{W}_{i},z^{S}_{i},i=1,\ldots,N=8760$ be actual time series data in an hourly resolution for $\log(W)$ and $S$, respectively. Realizations $y^{W}_{i}=z^{W}_{i}-\mu_{W}(i),i=1,\ldots,N$, of $Y^{W}$ are calculated by removing the seasonality $\mu_{W}$ from the data. The same holds for the realizations $y^{S}_{i}=z^{S}_{i}-\mu_{S}(i),i=1,\ldots,N$, of $Y^{S}$, which are calculated by removing the seasonality $\mu_{S}$. Regarding the autoregressive structure (153) and the fact that the random vector $(Y^{W},Y^{S})^{\top}$ is bivariate Gaussian, the conditional joint distribution is given by

### v3

@@ Appendix F Proofs
As for the proof above we use the definitions of $I_{W}(t),I_{S,1}(t)$ and $I_{S,2}(t)$. Noticing that all of these expressions are martingales with mean zero. Also $B^{W}$ and $B^{S}$ are independent Brownian motions and therefore $I_{W}(t)$ is independent of $I_{S,1}(t)$ and $I_{S,2}(t)$. Using the solutions for $Y^{W}(t)$ and $Y^{S}(t)$ in (98) and (106), respectively, leads to the following calculation for $\mathbb{E}[Y^{W}(t)Y^{S}(t)]$
$\mathbb{E}[Y^{W}(t)Y^{S}(t)]$ $=\mathbb{E}[(m_{Y^{W}}(\tau,y_{W})+I_{W}(t))(m_{Y^{S}}(\tau,y_{W},y_{S})+I_{S,1}(t)+I_{S,2}(t))]$
$=\mathbb{E}[m_{Y^{W}}(\tau,y_{W})m_{Y^{S}}(\tau,y_{W},y_{S})]+\mathbb{E}[m_{Y^{W}}(\tau,y_{W})I_{S,1}(t)]+\mathbb{E}[m_{Y^{W}}(\tau,y_{W})I_{S,2}(t)]$
$\hskip 8.5359pt+\mathbb{E}[I_{W}(t)m_{Y^{S}}(\tau,y_{W},y_{S})]+\mathbb{E}[I_{W}(t)I_{S,1}(t)]+\mathbb{E}[I_{W}(t)I_{S,2}(t)]$
$=m_{Y^{W}}(\tau,y_{W})m_{Y^{S}}(\tau,y_{W},y_{S})+\mathbb{E}[I_{W}(t)]\mathbb{E}[I_{S,1}(t)]+\mathbb{E}[I_{W}(t)I_{S,2}(t)]$
@@
Substitution into (116) and applying Itô isometry yield the following result for the covariance
$\Sigma_{WS}(\tau)$ $=\mathbb{E}[I_{W}(t)I_{S,2}(t)]$
$=\mathbb{E}\bigg[\bigg(\int_{t_{a}}^{t}\sigma_{W}e^{-\lambda_{W}(t-r)}\mathrm{d}B^{W}(r)\bigg)\bigg(-\tfrac{\lambda_{S}c_{W}}{\lambda_{S}-\lambda_{W}}\sigma_{W}\int_{t_{a}}^{t}(e^{-\lambda_{W}(t-r)}-e^{-\lambda_{S}(t-r)})\mathrm{d}B^{W}(r)\bigg)\bigg]$
$=\int_{t_{a}}^{t}-\tfrac{\lambda_{S}c_{W}}{\lambda_{S}-\lambda_{W}}\sigma_{W}^{2}e^{-\lambda_{W}(t-r)}(e^{-\lambda_{W}(t-r)}-e^{-\lambda_{S}(t-r)})\mathrm{d}r$
$=-\tfrac{\lambda_{S}c_{W}}{\lambda_{S}-\lambda_{W}}\sigma_{W}^{2}\int_{t_{a}}^{t}(e^{-2\lambda_{W}(t-r)}-e^{-(\lambda_{S}+\lambda_{W})(t-r)})\mathrm{d}r$
$=-\tfrac{\lambda_{S}c_{W}}{\lambda_{S}-\lambda_{W}}\sigma_{W}^{2}\bigg(\int_{t_{a}}^{t}e^{-2\lambda_{W}(t-r)}\mathrm{d}r-\int_{t_{a}}^{t}e^{-(\lambda_{S}+\lambda_{W})(t-r)}\mathrm{d}r\bigg)$
$=-\tfrac{\lambda_{S}c_{W}}{\lambda_{S}-\lambda_{W}}\sigma_{W}^{2}\bigg(\tfrac{\Sigma_{Y^{W}}^{2}(\tau)}{\sigma_{W}^{2}}-\tfrac{1}{(\lambda_{S}+\lambda_{W})}\bigg(1-e^{-(\lambda_{S}+\lambda_{W})\tau}\bigg)\bigg)$
@@
To show that $\Sigma_{WS}(\tau)\leq 0$, two different cases must be considered.
@@
In this case, it holds that $\lambda_{S}-\lambda_{W}>0$ and therefore $\tfrac{\lambda_{S}c_{W}}{\lambda_{S}-\lambda_{W}}\leq 0.$ Note that $\lambda_{S}+\lambda_{W}>2\lambda_{W}$ and using (100) this leads to
@@
It is $\lambda_{S}-\lambda_{W}<0$ and gives $\tfrac{\lambda_{S}c_{W}}{\lambda_{S}-\lambda_{W}}>0.$ Conversely, to the first case, we get $\lambda_{S}+\lambda_{W}<2\lambda_{W}$, leading to
@@
being zero if and only if $c_{W}=0$, which completes the proof.
@@
The matrix $\Sigma(\tau)$ is positive definite if $\operatorname{\mathrm{tr}\,}\Sigma(\tau)>0$ and $\det\Sigma(\tau)>0$. Because of $\Sigma_{W}^{2}(\tau),\Sigma_{S}^{2}(\tau)>0$, it is $\operatorname{\mathrm{tr}\,}\Sigma(\tau)=\Sigma_{W}^{2}(\tau)+\Sigma_{S}^{2}(\tau)>0$. Furthermore, $\det\Sigma(\tau)=\Sigma_{W}^{2}(\tau)\Sigma_{S}^{2}(\tau)-\Sigma_{WS}^{2}>0$ if $\Sigma_{W}^{2}(\tau)\Sigma_{S}^{2}(\tau)>\Sigma_{WS}^{2}(\tau)$. Let $\gamma=\bigg(\tfrac{\lambda_{S}c_{W}\sigma_{W}}{\lambda_{S}-\lambda_{W}}\bigg)^{2}$, then calculations yield
$\Sigma_{W}^{2}(\tau)\Sigma_{S}^{2}(\tau)$ $=\Sigma_{W}^{2}(\tau)\bigg[\Sigma_{Y^{S}}^{2}(\tau)+\gamma\big(\Sigma_{Y^{W}}^{2}(\tau)+\frac{\sigma_{W}^{2}}{\sigma_{S}^{2}}\Sigma_{Y^{S}}^{2}(\tau)-\frac{2\sigma_{W}^{2}}{\lambda_{S}+\lambda_{W}}\big(1-e^{-(\lambda_{S}+\lambda_{W})\tau}\big)\big)\bigg]$
$=\Sigma_{W}^{2}(\tau)\Sigma_{Y^{S}}^{2}(\tau)+\gamma\bigg(\Sigma_{Y^{W}}^{4}(\tau)+\frac{\sigma_{W}^{2}}{\sigma_{S}^{2}}\Sigma_{W}^{2}(\tau)\Sigma_{Y^{S}}^{2}(\tau)-\frac{2\Sigma_{W}^{2}(\tau)\sigma_{W}^{2}}{\lambda_{S}+\lambda_{W}}\big(1-e^{-(\lambda_{S}+\lambda_{W})\tau}\big)\bigg)$
$=\Sigma_{W}^{2}(\tau)\Sigma_{Y^{S}}^{2}(\tau)+\gamma\bigg(\bigg[\Sigma_{W}^{2}(\tau)-\frac{\sigma_{W}^{2}}{\lambda_{S}+\lambda_{W}}(1-e^{-(\lambda_{S}+\lambda_{W})\tau})\bigg]^{2}$
$\hskip 8.5359pt+\frac{\sigma_{W}^{2}}{\sigma_{S}^{2}}\Sigma_{W}^{2}(\tau)\Sigma_{Y^{S}}^{2}(\tau)-\frac{\sigma_{W}^{4}}{(\lambda_{S}+\lambda_{W})^{2}}(1-e^{-(\lambda_{S}+\lambda_{W})\tau})^{2}\bigg)$
@@
where $A=\frac{\sigma_{W}^{2}}{\sigma_{S}^{2}}\Sigma_{W}^{2}(\tau)\Sigma_{Y^{S}}^{2}(\tau)-\frac{\sigma_{W}^{4}}{(\lambda_{S}+\lambda_{W})^{2}}(1-e^{-(\lambda_{S}+\lambda_{W})\tau})^{2}$. Note that $\Sigma_{W}^{2}(\tau)\Sigma_{Y^{S}}^{2}(\tau)>0$ and $\gamma>0$, so we need to verify that $A\geq 0$ in order to show that $\Sigma$ is positive definite. Using (100) and (109), we obtain
@@
and by using $\int_{0}^{1}e^{-az}\,\mathrm{d}z=\frac{1-e^{-a}}{a}$ this leads to
@@
Applying the Cauchy-Schwarz inequality gives
@@
F.2 Proof of Proposition 3.2
Let $0\leq t_{a}<t_{b}\leq t_{\text{E}}$ with $t_{a}=t_{n}$ and $t_{b}=t_{n+1}$ and $(Y^{W}_{n},Y^{S}_{n})=(y_{W},y_{S})$. The pair $(\log W_{n+1},S_{n+1})$ is bivariate Gaussian, as it results from a linear transformation of the bivariate Gaussian $(Y^{W}_{n+1},Y^{S}_{n+1})$. Moreover, we have $\Delta t=\tau=t_{n+1}-t_{n}\equiv const$ and therefore the covariance matrix is given by $\Sigma=\Sigma_{Y}(\Delta t)=\Sigma_{Y}(\tau)$. The linear transformation leads to the following mean
@@
$\begin{split}m^{W}_{n+1}(w)&=\mu_{W}(t_{n+1})+m_{Y^{W}}(\Delta t,y_{W}),\\ m^{S}_{n+1}(w,s)&=\;\mu_{S}(t_{n+1})\,+m_{Y^{S}}(\Delta t,y_{W},y_{S}),\end{split}$ (128)
@@
F.3 Proof of Corollary 3.3
The Cholesky decomposition for $\Sigma$ can be verified by some simple calculations. Furthermore, recalling Proposition 3.2, the pair $(\log W,S)$ is bivariate Gaussian with mean $m_{WS}$ and covariance matrix $\Sigma$. Based on this, given a standard normally distributed random vector $Z_{n+1}=(Z^{W}_{n+1},Z^{S}_{n+1})^{\top}\sim\pazocal{N}(0_{2},I_{2})$, we can write $(\log W,S)$ as a linear transformation of the form
$(\log W_{n+1}S_{n+1})=m^{WS}_{n+1}(W_{n},S_{n})+AZ_{n+1},\quad\text{ with }A=\begin{pmatrix}\Sigma_{W}&0\\ \rho\Sigma_{S}&\sqrt{1-\rho^{2}}\Sigma_{S}\end{pmatrix}$. (129)
@@
Appendix G Calibration of Wind Speed and Energy Price Model
@@
The wind speed and the energy price are modeled by continuous-time Ornstein-Uhlenbeck processes on $0,t_{\text{E}}]$, see (16). Choosing appropriate parameters is important in order to obtain a realistic dynamical system. Therefore, real-world data is needed to calibrate the model parameters. Data for calibration can be used from the website of Deutscher Wetterdienst [^1] Deutscher Wetterdienst, Climate data center, https://opendata.dwd.de/climate_environment and ENTSO-E [^2] ENTSO-E, Transparency platform, https://transparency.entsoe.eu/dashboard/show. Note that the electricity price data is provided in a resolution of every 15 minutes or every hour. The latter results by averaging the 15 minute values. In our calibration, wind speed and electricity price are provided in an hourly resolution for the entire year 2020, i.e., 8760 data points each. The model assumes that data for wind speed is given by realizations of a random variable $e^{Z(t)}$ for a Gaussian process $Z$. Therefore, the fetched data has to be transformed by applying a logarithmic transformation. Denote the individual data points by $z^{\dagger}_{1},\ldots,z^{\dagger}_{8760}$ with $\dagger\in\{W,S\}$, where $z_{j}^{\dagger}$ is the data for the $j$-th hour for wind and price respectively. Moreover, outliers need to be removed from the dataset. This is realized by applying the $3$-$\sigma$ rule and remove values outside of the $99.7$ % confidence interval.
G.1 Estimation of the Seasonality Function
@@
For an approximation of the seasonality functions $\mu_{W}$ and $\mu_{S}$, an appropriate modeling function is fitted to this data by least-square regression. The following models where chosen in order capture the daily, half-daily and yearly trends in the underlying data: for wind speed
$\mu_{W}(t)=k^{W}_{0}+k^{W}_{1}\cos\bigg(\frac{2\pi(t-t^{W}_{1})}{8760}\bigg)+k^{W}_{2}\cos\bigg(\frac{2\pi(t-t^{W}_{2})}{24}\bigg),$ (130)
and electricity price
$\mu_{S}(t)=k^{S}_{0}+k^{S}_{1}\cos\bigg(\frac{2\pi(t-t^{S}_{1})}{8760}\bigg)+k^{S}_{2}\cos\bigg(\frac{2\pi(t-t^{S}_{2})}{24}\bigg)+k^{S}_{3}\cos\bigg(\frac{2\pi(t-t^{S}_{3})}{12}\bigg),$ (131)
@@
The denominator in the cosine functions reflects the daily, half-daily and yearly components of the data. The choice of these functions is motivated by the patterns in the data. Wind speeds are commonly higher in winter compared to summer. This fact is also reflected in the daily variations, where wind is slower at night and faster during the day. Electricity prices need to adjust to the changing weather conditions, especially with regard to renewable energy resulting in a yearly seasonality. This adjustments are also needed on a daily scale, arising due to day-to-night cycle of renewable energy sources. The price data reveals that there is a need to include a half-daily trend, due to the fact that most electricity peak load is typically used during two periods in a day: in the morning (people wake up) and in the afternoon (people return from work). The fitted parameters for the seasonalites can be found in Table 2.
@@
G.2 Estimation of Ornstein-Uhlenbeck Parameters
Recall that $\log W(t)=\mu_{W}(t)+Y^{W}(t)$ and $S(t)=\mu(t)+Y^{S}(t)$, where $Y^{W}$ and $Y^{S}$ in (17) are Ornstein-Uhlenbeck processes described in Lemma 3.1. The discrete-time analogue of the closed-form solution (18) is an autoregressive process of the form
$\begin{split}Y^{W}_{n+1}&=p_{W}Y_{n}^{W}+\Sigma_{Y^{W}}Z^{W}_{n+1},\\ Y^{S}_{n+1}&=p_{S}Y_{n}^{S}+q_{S}Y_{n}^{W}+\Sigma_{S}Z^{S}_{n+1},\end{split}$ (132)
with a sequence $(Z_{n})_{n=1,\ldots,N}$ of independent standard normally distributed random vectors $Z_{n+1}=(Z_{n+1}^{W},Z_{n+1}^{S})^{\top}\sim\pazocal{N}(0_{2},I_{2})$. The parameters $p_{W},p_{S},q_{S},\Sigma_{Y^{W}}^{2}$ and $\Sigma_{S}^{2}$ are identified with their associated quantities given in (20), specified as
$\begin{split}&p_{W}=e^{-\lambda_{W}\Delta t},\quad\Sigma_{Y^{W}}^{2}=\tfrac{\sigma_{W}^{2}}{2\lambda}(1-e^{-2\lambda_{W}\Delta t}),\\ &p_{S}=e^{-\lambda_{S}\Delta t},\quad q_{S}=-\frac{\lambda_{S}c_{W}}{\lambda_{S}-\lambda_{W}}(e^{-\lambda_{W}\Delta t}-e^{-\lambda_{S}\Delta t}),\\ &\Sigma_{S}^{2}=\Sigma_{Y^{S}}^{2}+\tfrac{(\lambda_{S}c_{W})^{2}}{(\lambda_{S}-\lambda_{W})^{2}}\bigg[\Sigma_{Y^{W}}^{2}+\tfrac{\sigma_{W}^{2}}{\sigma_{S}^{2}}\Sigma_{Y^{S}}^{2}-\tfrac{2\sigma_{W}^{2}}{\lambda_{S}+\lambda_{W}}(1-e^{-(\lambda_{S}+\lambda_{W})\Delta t})\bigg].\end{split}$ (133)
Let $z^{W}_{i},z^{S}_{i},~i=1,\ldots,N=8760$ be the actual time series data in an hourly resolution for $\log(W)$ and $S$, respectively. The realizations $y^{W}_{i}=z^{W}_{i}-\mu_{W}(i)$ of $Y^{W}$ are calculated by removing the seasonality $\mu_{W}$ from the data. The same holds for the realizations $y^{S}_{i}=z^{S}_{i}-\mu_{S}(i)$ of $Y^{S}$, which are calculated by removing the seasonality $\mu_{S}$. Regarding the autoregressive structure (132) and the fact that the random vector $(Y^{W},Y^{S})^{\top}$ is bivariate Gaussian, the conditional joint distribution is given by